### 17 Dec 2023 v3

@@ 4.3 Hypothetical interventions on the principal strata
Note, previous studies have shown a the link between reduced sBP and reduced all-cause mortality [2]. Hence, it is likely that an intervention targeting sBP not only may have an effect on cognition but also on mortality rates. In such situations, the $PCIE$ would compare potential outcomes for different underlying populations. In particular, if subjects live longer under the risk-reducing intervention this population may include less healthy subjects who would die under the natural course, masking a possible intervention effect. The $SAIE$ on the other hand, compare potential outcomes from the same population, i.e. a population consisting of subjects who would survive both when intervening on sBP and under the natural course. Thereby, providing a estimate that adjust for differences in mortality rates for the two regimes.
In our application we are interested in both a prognostic effect estimate (i.e. the $PCIE$) and an etiological effect estimate (i.e. the $SAIE$). For identification of the $SAIE$, we need two additional assumptions.
@@
C5 Stochastic Monotonicity: $\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]\geq\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=0]$. Given a beneficial intervention, we assume the probability of survival under the intervention $g^{int^{*}}$ is higher among those subjects that would survive under the natural course $g^{nc}$, compared to those who would not. We further let $\Pr[S_{T}(g)=1]=\psi^{g}$, for $g=(g^{nc},g^{int^{*}})$, and assume $\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]=\psi^{g^{int^{*}}}+\lambda\left[\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}-\psi^{g^{int^{*}}}\right]$, where $\lambda$ is a sensitivity parameter and $0\leq\lambda\leq 1$ [33, 21]. Note, if $\lambda=1$ this assumption corresponds to a deterministic monotonicity assumption where subjects who were to be alive under the observed regime $g^{nc}$ would also be alive under the intervention $g^{int^{*}}$, i.e. $S_{T}(g^{nc})\leq S_{T}(g^{int^{*}})$. In contrast, if $\lambda=0$ this corresponds to an assumption of $S_{T}(g^{nc})$ and $S_{T}(g^{int^{*}})$ being independent.
@@
C6 Difference in expectations of outcomes when comparing different strata. For $g\in(g^{int^{*}},g^{nc})$, $g^{\prime}\in(g^{int^{*}},g^{nc})$ and $g\neq g^{\prime}$, we assume, $E[Y_{T}(g)|S_{T}(g^{\prime})=S_{T}(g)=1]-E[Y_{T}(g)|S_{T}(g)=1,S_{T}(g^{\prime})\neq 1]=\Delta^{g}$ and further that $\Delta^{g}=\Delta^{g^{\prime}}=\Delta$. Note, if $\Delta=0$ there is no difference in potential outcomes when comparing the ”always survivor” strata to the strata where individuals were to live under regime $g$ but not under the contrasting regime $g^{\prime}$. In contrast, if $\Delta>0$ this implies a higher potential outcome is observed for the principal strata, i.e. the strata where individuals were to live under both regimes.
@@
A detailed description of identification of the $SAIE$ is given in Supplementary Appendix B. In short, under assumptions C1-C6, the $SAIE$ is given by:
$\quad PCIE+\Delta\left\{\psi^{g^{int^{*}}}+\lambda\left(U-\psi^{g^{int^{*}}}\right)\right\}\left(1-\frac{1}{U}\right),$ (4)
@@
where $U=\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}$. In particular, under A5 (stochastic monotonicity) (4) represents a location shift upwards from the $PCIE$ for a beneficial intervention where we expect the participants to improve their health under the intervention regime compared to the natural course. Note, $\psi^{g}=E_{H^{g}}\left\{\Pr[S_{T}=1|H^{g}]\right\}$, for $g\in(g^{int^{*}},g^{nc})$, are obtained by marginalizing over the distributions of the set of temporally preceding variables $H^{nc}$ using the g-formula and $H^{int^{*}}$ using the extended g-formula.
@@ 5 A semi-parametric approach for estimation
[…]
As for traditional (hard) BART models, we have that the mean function of a variable conditional on the history, is given by the sum of regression trees, $\mu_{\cdot}=\sum_{t=1}^{T}g_{y_{t}}(\mathcal{T}_{t},\mathcal{M}_{t})$, where $\mathcal{M}_{t}$ is the parameters for the leaf nodes and $\mathcal{T}_{t}$ the tree structure. In general, the distribution of the continuous variable are moreover specified as normal, and the binary variables are specified as probit models.
The regularization prior, controlling the complexity of the sum-of-trees, differ in two ways between the standard BART algorithm and soft BART. First, for soft BART, the path down the decision tree is probabilistic, where the sharpness of the decision is controlled by the bandwidth parameter. Thereby adapting to the smoothness level of the regression function even for low dimensional tree ensembles. Second, a predictor $x_{j}$, is chosen according to a probability vector $\mathbf{q}=(q_{1},\ldots,q_{P})$, typically specified as uniform. In contrast, soft BART models incorporate a Dirichlet distribution for $\mathbf{q}$, such that $\mathbf{q}\sim\mathcal{D}\left(\alpha/P,\ldots,\alpha/P\right)$. The sparsity-inducing prior was first introduced for high dimensional settings [22], and can be extended to allow sparsity both within and between groups [23, 6]. Determining the degree of sparsity in the model is done by placing a prior on $\alpha$, as such, $\frac{\alpha}{\alpha+\rho}\sim Beta\left(a,b\right)$. Default specifications for $a,b$ and $\rho$ is $0.5$, $1$ and $P$ respectively. This sparsity inducing prior is implemented for both hard and soft BART. We refer to BART models with this prior as Dirichlet additive regression trees (DART).
@@ 5.2 A grouping prior for longitudinal data
We consider a hierarchical structure where the predictors are divided into two groups: predictors measured at the current time point $t$ and predictors measured at past time points $k=1,\ldots,t-1$. Moreover, baseline predictors, e.g. age and sex, are grouped with the current predictors into one group, since we want to impose the same degree of sparsity on these predictors. Let $w$ refer to the probability of a predictor being from the set of current predictors, versus from the set of past predictors; we specify a beta prior, $w\sim Beta\left(a,b\right)$. For the current predictors, let $v^{t}_{j}$ refer to the probability of selecting the $j$th predictor from the set of current predictors $\textbf{v}^{t}=(v^{t}_{1},\ldots,v^{t}_{P_{t}}$); $\textbf{v}^{t}$ is given a Dirichlet prior $\textbf{v}^{t}\sim\mathcal{D}\left(\eta/P_{t},\ldots,\eta/P_{t}\right)$. So a priori, the probability of selecting the $j$th current predictor to construct a given split is equal to $w\times v^{t}_{j}$. The past predictors are divided into $t-1$ groups, where $P_{k}$ is the size of group associated with time $k$. Let $u^{k}$ denote the probability of selecting the $k$th set of time-varying covariates, and let $v^{k}_{j}$ denote the probability of selecting the $j$th predictor in the $k$th set of time-varying covariates. To introduce sparsity related to proximity in time to the response we give u a grouping Dirichlet prior $\textbf{u}\sim\mathcal{D}\left(\alpha_{1}/t-1,\ldots,\alpha_{t-1}/t-1\right)$, where we determine the degree of sparsity in the model by placing a prior on $\alpha_{k}$ such that a larger degree of sparsity is obtained for smaller $k$’s (see details below). Similarly $\textbf{v}^{k}$ is also given Dirichlet prior $\textbf{v}^{k}\sim\mathcal{D}\left(\varphi^{k}/P_{k},\ldots,\varphi^{k}/P_{k}\right)$. So a priori, the probability of selecting the $j$th predictor from the $k$th time point to construct a given split is equal to $(1-w)\times u^{k}\times v^{k}_{j}$.
For the hyperparameters $\eta$ and $\varphi^{k}_{j}$ we use the default Beta hyperprior as implemented in DART. For $\alpha_{j}$ we specify a Beta hyperprior $\frac{\alpha_{j}}{\alpha_{j}+\rho_{j}}\sim Beta\left(c_{j},d_{j}\right)$; we let $c_{j}$ vary as a linear function of time, set $d_{j}=1$, and $\rho_{j}=t$ for all $j$. To introduce sparsity related to proximity in time to the response we set $c_{j}=1-\frac{(t-j)}{t-1}*0.5$. This results in ordered hyperpriors for the $\alpha_{j}$s, where $c_{1}=0.5$ for predictors measured at the first time point and $10.5<c_{j-1}<c_{j}<$ 1 for $j>1$. That is, a higher degree of sparsity is introduced for predictors measured earlier in time and less sparsity for predictors measured closer in time to the response. The simulations in Section 6 shows the strong performance of these hyperprior choices. Details about the posterior computations of the L-Dirichlet prior are obtained in the Supplementary Appendix D.
@@ 5.3 Algorithm to compute the $PCIE$ and $SAIE$
The $PCIE$ and the $SAIE$ can be computed using the algorithm described in the Supplementary Appendix D and is implemented in the R package GcompLDART (available on https://github.com/m4ryjo/GcompLDART).
@@ 6 Simulation study
In this simulation study we evaluate the performance of L-DART and compare it to other BART specifications. Data were generated according to a modified version of the Friedman’s example [10], mirroring a longitudinal study with five time-varying predictors and a continuous outcome, with $f(\bar{\mathbf{x}}_{i})=\sum_{t=1}^{T}c_{t}[10\mathrm{sin}(\pi x_{t}^{1}x_{t}^{2})+20(x_{t}^{3}-0.5)^{2}+10x_{t}^{4}+5x_{t}^{5}]$, where $Y_{iT}\sim\mathrm{Normal}\left\{f(\bar{\mathbf{X}}_{i}),\sigma_{T}\right\}$, $\sigma_{T}=10$, and $c_{t}$ determines the strength of association between the predictors at time $t$ and the outcome.
@@
The longitudinal predictors, $\bar{\mathbf{X}}_{i}=\bar{X}_{i}^{1},\ldots,\bar{X}_{i}^{T}$, were generated as correlated uniform random variables, generated as multivariate normal variables, with zero mean and covariance as described below, and then transformed to uniform variables. As such, $\bar{X}_{it}^{p}\sim Uniform(0,1)$, where the between time-points covariance for predictor $X^{p}$ is $Var(X^{p}_{it},X^{p}_{ik})=d_{tk}^{2}$. Here, we specified $d_{jj-1}=0.4$, $d_{jj-2}=0.2$ and $d_{jj-3}=0.1$, reflecting a stronger correlation for measures closer in time, and all other were set to zero.
@@
We simulated 500 datasets of size $n=\left(100,500,1500\right)$, and five different scenarios by varying $\bar{c}=\left(c_{1},\ldots,c_{T}\right)$: Scenario A: $\bar{c}=(1,0.75,0.5,0.25)$, reflecting a decreasing, although a relatively strong, association over time between the predictors and the outcome; Scenario B: $\bar{c}=(1,0.75,0.5,0.25)^{3}$, i.e. a sharp decrease in association over time; Scenario C: $\bar{c}=(1,0,0,0)$, i.e. only predictors at the closest time point are associated with the outcome; Scenario D: $\bar{c}=(1,0.75,0.5,0.25,0,0,0,0)$, which is similar to Scenario A but also includes four time points uncorrelated with the outcome; and Scenario E: $\bar{c}=(0.25,0.25,0.25,0.25)$, i.e. a weak and stable association with the outcome for all time points. For scenarios A, B, C, and, E, $T=4$ and for scenario D, $T=8$. For each scenario, 10 or 100 trees were specified for each model.
We compared the predictive performance of our proposed approach (L-DART) to two hard BART models; the standard BART algorithm (H-BART) and BART with the sparsity-inducing Dirichlet prior (H-DART) (as implemented in the R package BART), and two soft BART models; the standard S-DART algorithm (as implemented in the R package SoftBart) and SoftBart incorporating only the predictors from the last time point in the model (S-DART-lag1). For evaluation of their performance, the methods were first trained on one dataset and compared on a test dataset using: % bias, root-mean-squared error (RMSE), and 95% coverage probabilities (CPs), for all the predictions of the test data. Computed as the posterior mean of: %$\%bias=\frac{1}{\sum_{i=1}^{n}\mu_{T}}\sum_{i=1}^{n}\frac{y_{iT}-\mu_{iT}}{n}$, where $\mu_{T}=\frac{1}{n}\sum_{i=1}^{n}\mu_{iT}$ and $\mu_{iT}$ are obtained from the training data; $RMSE=\sqrt{\sum_{i=1}^{n}\frac{(y_{iT}-\mu_{iT})^{2}}{n}}$; and the $95\%$ CP is the average pointwise predictive coverage of $95\%$ CIs for all the predictions of the test data set. The models use their respective default priors (if not stated otherwise) and were fitted by using 3000 warm-up iterations, 400000 sampling iterations, and a thinning of 100. All simulations were done using R-4.1 at the High Performance Computing Center North (HPC2N) and code is available on https://github.com/m4ryjo/GcompLDART.

### 20 Mar 2024 v4

@@ 4.3 Inference on the always survivor stratum
Conditioning on survival as described in Section 4.1 may introduce bias due to the fact that survival is a post-randomization event [41]. Hence, the partly conditional approach in Section 4.2 cannot address etiological research questions in mortal cohorts. An alternative estimand is the causal effect on the subpopulation who would survive through the end of follow-up regardless of treatment regime ([6, 7]). Our goal here is to estimate the difference in expected potential outcomes between $g^{nc}$ and $g^{int^{*}}$ in the always survivor stratum. That is, what we call a survivors average intervention effect: $SAIE=E[Y(g^{int^{*}})-Y(g^{nc})\mid S_{T}(g^{nc})=S_{T}(g^{int^{*}})=1]$. Note, previous studies have shown a the link between reduced sBP and reduced all-cause mortality [1]. Hence, it is likely that an intervention targeting sBP not only may have an effect on cognition but also on mortality rates. In such situations, the $PCIE$ would compare potential outcomes for different underlying populations. In particular, if subjects live longer under the risk-reducing intervention this population may include less healthy subjects who would die under the natural course, masking a possible intervention effect. The $SAIE$ on the other hand, compare potential outcomes from the same population, i.e. a population consisting of subjects who would survive both when intervening on sBP and under the natural course. As such, it provides an estimate that adjusts for differences in mortality rates between the two regimes. In our application we are interested in both a prognostic effect estimate (i.e. the $PCIE$) and an etiological effect estimate (i.e. the $SAIE$).
For identification of the $SAIE$, we need two additional assumptions.
@@
C5 Stochastic Monotonicity: $\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]\geq\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=0]$. In the current application, we impose an intervention that improves individuals’ health. As such, we assume the probability of survival under the intervention $g^{int^{*}}$ is higher among those subjects that would survive under the natural course $g^{nc}$, compared to those who would not. We denote $\Pr[S_{T}(g)=1]$ by $\psi^{g}$ for $g$ in $\{g^{nc},g^{int^{*}}\}$, and we assume $\psi^{g^{int^{*}}}\geq\psi^{g^{nc}}$. We further assume $\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]=\psi^{g^{int^{*}}}+\lambda\left[\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}-\psi^{g^{int^{*}}}\right]$, where $\lambda$ is a sensitivity parameter [29, 18]. Note, $0\leq\lambda\leq 1$. If $\lambda=1$ this assumption corresponds to a deterministic monotonicity assumption where subjects who were to be alive under the observed regime $g^{nc}$ would also be alive under the intervention $g^{int^{*}}$, i.e. $S_{T}(g^{nc})\leq S_{T}(g^{int^{*}})$. In contrast, if $\lambda=0$ this corresponds to an assumption of $S_{T}(g^{nc})$ and $S_{T}(g^{int^{*}})$ being independent, which we in the current application expect to be unlikely.
@@
C6 Difference in expectations of outcomes when comparing different strata. For $g\in\{g^{int^{*}},g^{nc}\}$, $g^{\prime}\in\{g^{int^{*}},g^{nc}\}$ and $g\neq g^{\prime}$, we assume, $E[Y_{T}(g)|S_{T}(g^{\prime})=S_{T}(g)=1]-E[Y_{T}(g)|S_{T}(g)=1,S_{T}(g^{\prime})\neq 1]=\Delta^{g}$, where $\Delta^{g}$ is a sensitivity parameter. We further assume that $\Delta^{g}=\Delta^{g^{\prime}}=\Delta$, i.e. that the difference in expectations of potential outcomes when comparing different strata is the same for the two contrasting regimes. Note, if $\Delta=0$ there is no difference in potential outcomes when comparing the always survivor strata to the strata where individuals were to live under regime $g$ but not under the contrasting regime $g^{\prime}$. In contrast, if $\Delta>0$ this implies a higher potential outcome is observed for the always survivor stratum. In the current application we expect higher memory scores, i.e. a better memory, for the always survivor stratum.
@@
The identification details of the $SAIE$ are given in Supplementary Appendix B. Briefly, under assumptions C1-C6, the $SAIE$ is given by:
@@
$PCIE+\Delta\left\{\psi^{g^{int^{*}}}+\lambda\left(U-\psi^{g^{int^{*}}}\right)\right\}\left(1-\frac{1}{U}\right)$, where $U=\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}$. The equation represents a location shift upwards from the $PCIE$ for a beneficial intervention where we expect the participants to improve their health under the intervention compared to the natural course. Note, $\psi^{g}=E_{H^{g}}\left\{\Pr[S_{T}=1|H^{g}]\right\}$, for $g\in\{g^{int^{*}},g^{nc}\}$, are computed by marginalizing over the distributions of the set of temporally preceding variables $H^{nc}$ using the g-formula and $H^{int^{*}}$ using the extended g-formula.
@@ 5 A semi-parametric approach for estimation
In this section, we develop a Bayesian semi-parametric (BSP) modeling approach using soft Bayesian additive regression trees (BART) models to estimate each of the conditional distributions in (S0.Ex1) and (S0.Ex3) in Supplementary appendix A. The soft BART model improves upon the traditional hard BART model by addressing the lack of smoothness inherent in ensemble approaches. To enhance valid inference, especially in scenarios with multiple time points and a large set of regressors, we extend the current soft BART algorithm by introducing a novel sparsity-inducing hyperprior for longitudinal data. This longitudinal Dirichlet prior promotes parsimony by penalizing groups of predictors less likely to be important in longitudinal prediction settings. We refer to our model as a longitudinal Dirichlet additive regression trees (LDART) approach.
As for traditional (hard) BART models, we have that the mean function of a variable conditional on the history, is given by the sum of regression trees, $\mu_{t}=\sum_{b=1}^{B}g_{y_{t}}(\mathcal{T}_{b},\mathcal{M}_{b})$, where $\mathcal{M}_{b}$ is the parameters for the leaf nodes and $\mathcal{T}_{b}$ the tree structure. In general, the distribution of the continuous variable are moreover specified as normal, and the binary variables are specified as probit models.
The regularization prior, controlling the complexity of the sum-of-trees, differ in two ways between the standard BART algorithm and soft BART. First, for soft BART, the path down the decision tree is probabilistic, where the sharpness of the decision is controlled by the bandwidth parameter. Thereby adapting to the smoothness level of the regression function even for low dimensional tree ensembles, a known shortcoming for hard) BART models. Second, the $j$th predictor, from the set of $P$ predictors, is chosen according to a probability vector $\mathbf{q}=(q_{1},\ldots,q_{P})$, typically specified as uniform. In contrast, soft BART models incorporate a Dirichlet distribution for $\mathbf{q}$, such that $\mathbf{q}\sim\mathcal{D}\left(\alpha/P,\ldots,\alpha/P\right)$. The sparsity-inducing prior was first introduced for high dimensional settings [19], and can be extended to allow sparsity both within and between groups [20, 3]. Determining the degree of sparsity in the model is done by placing a prior on $\alpha$, as such, $\frac{\alpha}{\alpha+\rho}\sim Beta\left(a,b\right)$. Default specifications for $a,b$ and $\rho$ is $0.5$, $1$ and $P$ respectively. This sparsity inducing prior is implemented for both hard and soft BART. We refer to BART models with this prior as Dirichlet additive regression trees (DART).
@@ 5.2 A grouping prior for longitudinal data
We consider a hierarchical structure where the predictors are divided into two groups: predictors measured at the current time point $t$ and predictors measured at past time points $k=1,\ldots,t-1$. Moreover, baseline predictors, e.g. age and sex, are grouped with the current predictors into one group, since we want to impose the same degree of sparsity on these predictors. Let $w$ refer to the probability of a predictor being from the set of current predictors, versus from the set of past predictors; we specify a beta prior, $w\sim Beta\left(a,b\right)$. For the current predictors, let $v^{t}_{j}$ refer to the probability of selecting the $j$th predictor from the set of $P_{t}$ current predictors $\textbf{v}^{t}=(v^{t}_{1},\ldots,v^{t}_{P_{t}}$); $\textbf{v}^{t}$ is given a Dirichlet prior $\textbf{v}^{t}\sim\mathcal{D}\left(\eta/P_{t},\ldots,\eta/P_{t}\right)$. So a priori, the probability of selecting the $j$th current predictor to construct a given split is equal to $w\times v^{t}_{j}$. The past predictors are divided into $t-1$ groups, where $P_{k}$ is the size of group associated with time $k$. Let $u^{k}$ denote the probability of selecting the $k$th set of time-varying predictors, and let $v^{k}_{j}$ denote the probability of selecting the $j$th predictor in the $k$th set of time-varying predictors. To introduce sparsity related to proximity in time to the response we give u a grouping Dirichlet prior $\textbf{u}\sim\mathcal{D}\left(\alpha_{1}/t-1,\ldots,\alpha_{t-1}/t-1\right)$, where we determine the degree of sparsity in the model by placing a prior on $\alpha_{k}$ such that a larger degree of sparsity is obtained for smaller $k$’s (see details below). Similarly $\textbf{v}^{k}$ is also given Dirichlet prior $\textbf{v}^{k}\sim\mathcal{D}\left(\varphi^{k}/P_{k},\ldots,\varphi^{k}/P_{k}\right)$. So a priori, the probability of selecting the $j$th predictor from the $k$th time point to construct a given split is equal to $(1-w)\times u^{k}\times v^{k}_{j}$.
For the hyperparameters $\eta$ and $\varphi^{k}$ we use the default Beta hyperprior as implemented in DART. For $\alpha_{j}$ we specify a Beta hyperprior $\frac{\alpha_{j}}{\alpha_{j}+\rho_{j}}\sim Beta\left(c_{j},d_{j}\right)$; we let $c_{j}$ vary as a linear function of time, set $d_{j}=1$, and $\rho_{j}=t$ for all $j$. To introduce sparsity related to proximity in time to the response we set $c_{j}=1-\frac{(t-j)}{t-1}\times 0.5$. This results in ordered hyperpriors for the $\alpha_{j}$s, where $c_{1}=0.5$ for predictors measured at the first time point and $10.5<c_{j-1}<c_{j}<$ 1 for $j>1$. That is, a higher degree of sparsity is introduced for predictors measured earlier in time and less sparsity for predictors measured closer in time to the response. The simulations in Section 6 shows the strong performance of these hyperprior choices. Details about the posterior computations of the L-Dirichlet (LDART) prior are obtained in the Supplementary Appendix C.
@@ 5.3 Algorithm to compute the $PCIE$ and $SAIE$
The $PCIE$ and the $SAIE$ can be computed using the algorithm described in the Supplementary Appendix C and is implemented in the R package GcompLDART (available upon request).
@@ 6 Simulation study
In this simulation study we evaluate the performance of LDART and compare it to other BART specifications. Data were generated according to a modified version of the Friedman’s example [8], mirroring a longitudinal study with five time-varying predictors and a continuous outcome, with $f(\bar{\mathbf{x}}_{i})=\sum_{t=1}^{T}c_{t}[10\mathrm{sin}(\pi x_{t}^{1}x_{t}^{2})+20(x_{t}^{3}-0.5)^{2}+10x_{t}^{4}+5x_{t}^{5}]$, where $Y_{iT}\sim\mathrm{Normal}\left\{f(\bar{\mathbf{X}}_{i}),\sigma_{T}\right\}$, $T=4$, $\sigma_{T}=10$, and $c_{t}$ determines the strength of association between the predictors at time $t$ and the outcome.
@@
The longitudinal predictors, $\bar{\mathbf{X}}_{i}=\bar{\mathbf{X}}_{i}^{1},\ldots,\bar{\mathbf{X}}_{i}^{T}$, were generated as correlated uniform random variables, generated as multivariate normal variables, with zero mean and covariance as described below, and then transformed to uniform variables. As such, $\bar{X}_{it}^{p}\sim Uniform(0,1)$, where the between time-points covariance for predictor $X^{p}$ is $Var(X^{p}_{it},X^{p}_{ik})=d_{tk}^{2}$. Here, we specified $d_{jj-1}=0.4$, $d_{jj-2}=0.2$ and $d_{jj-3}=0.1$, reflecting a stronger correlation for measures closer in time. All correlations between different predictors were set to zero. We consider missingness in the form of dropout and assumed approximately $15\%$ dropped out at each follow-up time point $t=2,\ldots,T$. As such, dropout at time $j$ was generated as, $\pi_{j}(\bar{\mathbf{x}}_{i})=h_{j}\sum_{t=1}^{j-1}[x_{it}^{1}+x_{it}^{2}+x_{it}^{3}+x_{it}^{4}+x_{it}^{5}]$, where $R_{ij}\sim\mathrm{Binomial}\left\{\pi_{j}(\bar{\mathbf{x}}_{i})\right\}$ and $\bar{h}=(0.75,0.4,0.25)$.
@@
We simulated 650 datasets of size $n=\left(250,750,1500\right)$, and four different scenarios by varying $\bar{c}=\left(c_{T},\ldots,c_{1}\right)$: Scenario A: $\bar{c}=(1,0.75,0.5,0.25)$, reflecting a decreasing, although a relatively strong, association over time between the predictors and the outcome; Scenario B: $\bar{c}=(1,0.75,0.5,0.25)^{3}$, i.e. a sharp decrease in association over time; Scenario C: $\bar{c}=(1,0,0,0)$, i.e. only predictors at the closest time point are associated with the outcome; Scenario D: $\bar{c}=(0.25,0.25,0.25,0.25)$, i.e. a weak and stable association with the outcome for all time points. For all scenarios we estimated $E_{H^{g^{nc}}}\{E[Y_{T}(g^{nc})\mid H^{g^{nc}}]\}$ using equation (S0.Ex1).
We compared the performance of our proposed approach (LDART) with two hard BART models: the standard BART algorithm (HBART) and BART with the sparsity-inducing Dirichlet prior (HDART) (as implemented in the R package BART), as well as two soft BART models: the standard SoftBart algorithm (SDART; as implemented in the R package SoftBart) and SoftBart incorporating only the predictors from the last time point in the model (SDART-lag1). We specified 50 regression trees for each model. To evaluate their performance, we compared relative absolute bias, root-mean-squared error (RMSE), and 95% coverage probabilities (CPs).
The models use their respective default priors (if not stated otherwise) and used 1200 warm-up iterations, 2000 sampling iterations, a thinning of 4, and 48 chains. All simulations were done using R-4.1 at the High Performance Computing Center North (HPC2N) and code is available from the author upon request.